2) A sample of ?=24 observations is drawn from a normal population with ?=1000 and ?=240. Find each of the following:

A. ?(?¯>1097)

Probability =

B. ?(?¯<906)

Probability =

C. ?(?¯>990)

Probability =
Answer #1

sample size (n) = 24 < 30...but the sample is from a normal population..so, the sampling distribution of the sample mean will follow normal distribution with :-

mean (\mu_{\bar{x}})=\mu=1000

sd (\sigma_{\bar{x}})=\frac{\sigma}{\sqrt{n}}=\frac{240}{\sqrt{24}}=48.9898

a).\mathbf{P(\bar{X}>1097)}

=P(\frac{\bar{X}-1000}{48.9898}>\frac{1097-1000}{48.9898})

=P(Z>1.98)

=1-P(Z<1.98)

=1-\Phi(1.98)

=1-0.9761[ using standard normal table ]

=\mathbf{0.0239}

b).\mathbf{P(\bar{X}<906)}

=P(\frac{\bar{X}-1000}{48.9898}<\frac{906-1000}{48.9898})

=P(Z<-1.9188)

=\Phi(-1.9188)

=\mathbf{0.0275} [ in any blank cell of excel type =NORMSDIST(-1.9188) ]

c).\mathbf{P(\bar{X}>990)}

=P(\frac{\bar{X}-1000}{48.9898}>\frac{990-1000}{48.9898})

=P(Z>-0.2041)

=1-P(Z<-0.2041)

=1-\Phi(-0.2041)

=1-0.4191

=\mathbf{0.5809} [ in any blank cell of excel type =NORMSDIST(-0.2041) ]

*** if you have any doubt regarding the problem ,please write it in the comment box...if satisfied,please UPVOTE.